Two years ago, Anton was three times as old as his sister, Nathalie.In 4 years,He will be only twice his sister's age.Find their present ages.
Answer by richwmiller(17219)

Anton's present age=a
Nathalie's present age=n
a-2=3*(n-2)
a-2=3n-6
add 2 to both sides
a=3n-4
a+4=2*(n+4)
substitute 3n-4 for a
(3n-4)+4=2*(n+4)